OpenStudy (anonymous):

The difference of a number squared and the number is one less than four times the number

OpenStudy (triciaal):

let the number = x x^2 - 1 = 4 x-1 do you need to find the number x? x^2 -4 x = 0 x(x-4) = 0 x = 0; x- 4 = 0, x = 4 check x = 0 0^2-1 =-1 4(0) -1 = -1 x = 4 4^2 - 1 = 16-1 = 15 4*4 -1 = 16 - 1 = 15